OpenStudy (anonymous):

Matilda practices guitar for a total of 70 minutes every day. She spends 10 minutes more practicing the electric guitar than the acoustic guitar. Part A: Write a pair of linear equations to show the relationship between the number of minutes Matilda practices the acoustic guitar every day (x) and the number of minutes she practices the electric guitar every day (y). Part B: How much time does Matilda spend on practicing the acoustic guitar every day? Part C: Is it possible for Matilda to have spent 50 minutes practicing the electric guitar every day? Explain your reason?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@ertdre

OpenStudy (a1234):

So basically you have the equations 70 = x + y and y = x + 10.

OpenStudy (anonymous):

OK now what do I do?

OpenStudy (a1234):

Can you solve the system of equations?

OpenStudy (anonymous):

Kinda.

OpenStudy (a1234):

Give it a try.

OpenStudy (anonymous):

ok give a min

OpenStudy (anonymous):

\[x + y = 70 \]\[y = x + 10\] \[x + x + 10 = 70\]\[x = 30\] \[30 + y = 70\] \[y = 40\]

OpenStudy (a1234):

Right!

OpenStudy (a1234):

So the acoustic guitar is represented by x, so she spends 30 mins on it every day.

OpenStudy (anonymous):

so then she spends time on the electric guitar for 40 min

OpenStudy (a1234):

Yes

OpenStudy (anonymous):

but what do I do now?

OpenStudy (a1234):

You just solved Parts 1 and 2.

OpenStudy (anonymous):

I did? cool so for part c its no? @a1234

OpenStudy (a1234):

Yes. If you want to give your teacher a more detailed explanation, plug in 50 for y and show that it doesn't work.